### 10 Apr 2026 v1

# Tora3: Trajectory-Guided Audio-Video Generation with Physical Coherence

Abstract
Audio-video (AV) generation has recently made strong progress in perceptual quality and multimodal coherence, yet generating content with plausible motion-sound relations remains challenging. Existing methods often produce object motions that are visually unstable and sounds that are only loosely aligned with salient motion or contact events, largely because they lack an explicit motion-aware structure shared by video and audio generation. We present Tora3, a trajectory-guided AV generation framework that improves physical coherence by using object trajectories as a shared kinematic prior. Rather than treating trajectories as a video-only control signal, Tora3 uses them to jointly guide visual motion and acoustic events. Specifically, we design a trajectory-aligned motion representation for video, a kinematic-audio alignment module driven by trajectory-derived second-order kinematic states, and a hybrid flow matching scheme that preserves trajectory fidelity in trajectory-conditioned regions while maintaining local coherence elsewhere. We further curate PAV, a large-scale AV dataset emphasizing motion-relevant patterns with automatically extracted motion annotations. Extensive experiments show that Tora3 improves motion realism, motion-sound synchronization, and overall AV generation quality over strong open-source baselines.
Figure 1: Examples generated by Tora3. Tora3 generates audio-video content that follows prescribed trajectories with improved physical coherence, producing more plausible motion and better motion-sound alignment. Long trajectories are segmented for clarity. Please refer to the supplementary video for more examples.
Introduction
Recent advances in generative modeling have substantially improved video synthesis across multiple modalities, including text-to-video (T2V) [46, 3, 25, 62], audio-to-video (A2V) [8, 40, 49], and video-to-audio (V2A) [7, 16, 15]. More recently, unified text-to-audio-video (T2AV) models [19, 31, 34, 10, 12] have emerged, aiming to jointly synthesize visual and acoustic content from a single prompt. Commercial systems such as Veo3 [12], Sora2 [34], and Wan2.5 [10] suggest that high-fidelity multimodal generation is becoming increasingly practical.
Despite this progress, current T2AV generation is still driven primarily by semantic alignment rather than explicit motion-sound modeling. As a result, even strong open-source models [19, 31, 47] often produce visually implausible motion or audio that is only weakly synchronized with salient physical events [21]. In the visual domain, generated objects may follow inconsistent trajectories, exhibit implausible falling patterns, or undergo abrupt accelerations that break temporal continuity [13, 5]. In the acoustic domain, generated sound is often detached from the underlying motion. Impact sounds may miss the moment of contact, sound intensity may fail to reflect interaction strength, and continuous sounds such as sliding or driving often do not evolve with changes in motion. These failures are especially noticeable in interaction-heavy scenes and reduce realism. Existing T2AV methods typically align audio and video at a high semantic level, but lack an explicit intermediate representation that constrains both modalities under a shared motion-aware structure.
In this work, we focus on a practical notion of physical coherence in audio-video generation: whether generated object motion follows plausible trajectory-level dynamics, whether salient sound events are temporally aligned with motion or contact, and whether audio intensity evolves coherently with motion strength. We argue that object trajectories provide an effective and practical interface for introducing such structure into audio-video generation. Trajectories compactly describe where objects are and how they move over time, while exposing kinematic cues such as event timing, motion intensity, and temporal evolution. From this perspective, trajectories offer a natural shared representation for both modalities. They constrain plausible motion in video and indicate when sounds should occur and how their amplitude and temporal structure should evolve. Although trajectories do not fully specify real-world physics, they provide a lightweight and interpretable prior for improving motion-sound coherence in multimodal generation. Figure 2 shows that trajectory guidance better aligns visual motion with corresponding audio events, improving cross-modal synchronization.
[…]
Motivated by these observations, we present Tora3, a trajectory-guided audio-video generation framework that improves physical coherence through shared kinematic conditioning (see Figure 1 for examples). Rather than using trajectories as an auxiliary control signal, Tora3 treats them as a shared kinematic prior to jointly guide video motion and audio generation. Tora3 has three key components. First, for the video branch, we propose an efficient trajectory-aligned motion representation that reuses first-frame latents along sparse trajectories to inject motion cues directly into the latent space without an explicit motion encoder. This design avoids unnecessary distribution shift and preserves consistent motion conditioning throughout the AV pipeline. Second, for the audio branch, we design a kinematic-audio alignment module. Inspired by second-order motion representations [30, 60], we derive compact kinematic descriptors, including position, velocity, acceleration, and their magnitudes, from trajectories and inject them into the audio diffusion model via cross-attention. These trajectory-derived kinematic states provide direct cues for event timing and motion intensity, enabling sound generation whose loudness and temporal evolution better follow the underlying motion. Third, we introduce a hybrid flow matching mechanism with separate probability flows for trajectory and background regions. This mechanism preserves faithful trajectory-guided motion during early denoising while maintaining local appearance coherence during later refinement. To support training and evaluation, we further curate PAV, a large-scale audio-video dataset containing 460k video clips that emphasize motion-relevant patterns and provide automatically extracted motion annotations for trajectory-aware and kinematics-aware AV generation.
Our main contributions are as follows:
  • We propose Tora3, a trajectory-guided audio-video generation framework that uses object trajectories as a shared kinematic prior to jointly guide visual motion and acoustic events, improving physical coherence in generated AV content.
  • We develop a unified trajectory-grounded generation framework with three key components: (1) a trajectory-aligned motion representation for direct video latent conditioning, (2) a kinematic-audio alignment module based on trajectory-derived position, velocity, and acceleration cues, and (3) a hybrid flow matching scheme that preserves motion fidelity while maintaining local consistency.
  • We build PAV, a 460k-clip motion-centric AV dataset with automatic motion annotations, and show through extensive experiments that Tora3 consistently outperforms competitive baselines in video quality, trajectory faithfulness, and motion-sound coherence.
Related work
Trajectory-Controlled Video Generation
Trajectory-controlled video generation has attracted increasing attention for enabling fine-grained manipulation of object motion and scene dynamics. Prior works [28, 41, 59, 51, 27, 63, 9] inject motion cues such as optical flow [59, 41], point tracks [51, 63, 64], or temporally varying bounding boxes [11, 27] into diffusion models via dedicated encoders, adapters, or attention-based fusion modules. To better represent complex motions, some methods [54, 48] further explore 3D trajectories or depth-augmented keypoint maps, while others [17, 9] directly use pretrained base models without architectural modifications to improve scalability. Overall, these approaches substantially improve visual motion controllability and realism under trajectory guidance. However, existing trajectory-controlled methods mainly target video-only generation and do not address joint audio-video synthesis. In contrast, Tora3 extends trajectory guidance to unified audio-video generation by using trajectories as a shared kinematic prior for both motion control and audio-motion alignment.
Audio-Video Generation
Most prior unified audio-video generation methods [29, 39, 50, 56, 61] synthesize ambient sounds using dual-branch architectures, in which audio and video streams are modeled separately and fused through cross-attention or feature concatenation. Large-scale systems such as Veo3 show strong text-conditioned joint generation quality, and recent open-source models such as UniVerse-1 [47], Ovi [31], and LTX-2 [19] largely follow the same dual-backbone fusion paradigm. JoVA [20] further enables direct cross-modal interaction for human-centric content. Nevertheless, these methods often struggle to generate dynamic sequences in which audio remains well aligned with object motion and interaction timing. In contrast, Tora3 introduces trajectory-derived kinematic conditioning into joint AV generation. Specifically, we use object trajectories as a shared kinematic prior, combining trajectory-guided motion control in the video branch with kinematic-audio alignment in the audio branch. This design improves motion-sound coherence by aligning sound events and intensity changes with trajectory-level motion dynamics.
Figure 3: The overall architecture of Tora3. Object trajectories provide a shared kinematic prior for both video and audio generation. The video branch uses trajectory-aligned motion injection, while the audio branch incorporates trajectory-derived kinematic features to better align sound with motion. Hybrid flow matching further improves trajectory-guided generation by balancing motion fidelity and local coherence.
Methods
Overview
Tora3 is an audio-video generation framework built on a twin diffusion transformer architecture with separate video and audio DiT backbones, following Ovi [31]. It uses object trajectories as a shared kinematic prior across both modalities and consists of three components: trajectory-aligned motion representation for video, kinematic-audio alignment module driven by second-order kinematic states for audio, and hybrid flow matching for preserving motion fidelity and local coherence. The overall framework is illustrated in Figure 3.
Trajectory-aligned Motion Representation
Following Tora [63] and Tora2 [64], we use object trajectories to represent both local motion cues and longer-range temporal evolution. A common strategy in controllable video generation is to first encode trajectories with an auxiliary motion encoder and then inject the resulting features into the generative backbone. However, directly adopting such a design in a unified audio-video framework is suboptimal. Since both audio and video branches are large DiT backbones, introducing an additional motion encoder would further entangle optimization and make it harder to preserve a stable motion representation across modalities. In addition, extra encoding layers may attenuate sparse trajectory signals before they reach the backbone, which can reduce motion fidelity.
To avoid these issues, Tora3 directly maps trajectories into the latent space of the video backbone without introducing a dedicated motion encoder. We follow a standard image-to-video pipeline. Let $z\in\mathbb{R}^{h\times w\times d}$ denote the latent of the first frame, obtained by a pretrained VAE encoder [24], where $h$ and $w$ are the latent spatial dimensions and $d$ is the channel dimension. Let $x_{0}\in\mathbb{R}^{f\times h\times w\times d}$ denote the clean video latent sequence, where $f$ is the number of latent frames and the first frame corresponds to the conditioning image. Let $\widetilde{\mathcal{T}}$ denote the full object trajectories in pixel space. We apply temporal average pooling to $\widetilde{\mathcal{T}}$ to match the sampled video frames, obtaining $\mathcal{T}\in\mathbb{R}^{f\times N\times 2}$. We further map $\mathcal{T}$ to latent coordinates by scaling with the VAE downsampling factor and applying nearest-neighbor rounding, yielding $\mathcal{T}^{\prime}$, where each $\mathcal{T}^{\prime}_{i,n}=(p_{i,n},q_{i,n})\in\{0,\dots,h-1\}\times\{0,\dots,w-1\}$. Here, $p_{i,n}$ and $q_{i,n}$ denote the row and column index, respectively, of object $n$ at frame $i$ on the latent grid. We construct a trajectory-conditioned latent $x^{\mathrm{traj}}\in\mathbb{R}^{f\times h\times w\times d}$ by propagating first-frame object features along the trajectories. Specifically, for each object trajectory, we replace the latent feature at the trajectory-aligned location in frame $i$ with the corresponding first-frame feature:
$x^{\mathrm{traj}}[i,p_{i,n},q_{i,n},:]=z[p_{0,n},q_{0,n},:],\quad i\in\{1,\dots,f-1\}.$ (1)
If multiple trajectories map to the same latent location, we randomly select one. All non-trajectory locations in $x^{\mathrm{traj}}$ remain zero, except in the first frame, where they are overwritten by $z$. This design is motivated by two empirical observations: First, VAE latents are spatially smooth, so nearby spatial locations tend to share similar local appearance statistics [38]. Second, relocating latent features to nearby trajectory-aligned positions often preserves local object cues [52]. Together, these observations make first-frame latent propagation a direct and effective way to inject motion cues into the native latent space. This strategy provides an explicit trajectory-aligned conditioning signal while avoiding the additional feature transformation introduced by a separate motion encoder.
Kinematic-Audio Alignment
Semantic conditioning alone is insufficient for realistic motion-dependent audio synchronization. In monaural audio generation, motion-sound alignment depends heavily on temporal structure, spectral texture, and amplitude evolution, all of which are often correlated with object motion. Without explicit kinematic cues, the model must infer these relations implicitly from visual features, often leading to inaccurate timing and mismatched intensity.
Kinematic Feature Representation. To introduce explicit motion-aware structure into audio generation, we condition the audio branch on compact kinematic features derived from object trajectories. These features form a second-order kinematic state, since an object’s motion can be naturally described by its position, velocity, and acceleration. Together, they provide informative cues about when interactions occur, how strong they are, and how acoustic energy may evolve over time.
Given temporal-downsampled trajectories $\mathcal{T}\in\mathbb{R}^{f\times N\times 2}$, we normalize the 2D coordinates by the longer side of the image and denote the normalized position of object $n$ at frame $i$ as $\bm{r}_{i,n}\in[0,1]^{2}$. These normalized positions provide a coarse spatial anchor and useful scene context, such as whether an object is close to the ground or suspended in the air, which may help infer the underlying interaction type and associated sound.
We estimate velocity and acceleration using second-order central differences for interior frames, with forward and backward differences at the boundaries. Let $\tau=1/\text{fps}$ denote the time interval between consecutive frames. The velocity and acceleration at frame $i$ for trajectory point $n$ are approximated as
$\bm{v}_{i,n}\approx\frac{\bm{r}_{i+1,n}-\bm{r}_{i-1,n}}{2\tau},$ (2)
$\bm{a}_{i,n}\approx\frac{\bm{r}_{i+1,n}-2\bm{r}_{i,n}+\bm{r}_{i-1,n}}{\tau^{2}},$ (3)
valid for interior frames $i\in\{1,\dots,f-2\}$. This provides stable, second-order accurate estimates of trajectory-level motion. We explicitly retain both vector components and magnitudes to disentangle direction from strength.
The velocity vector characterizes the mode of interaction, for example, predominantly vertical motion for falling and impact, or horizontal motion for sliding and rolling, which can be related to different sound patterns. Its magnitude $\|\bm{v}_{i,n}\|_{2}$ reflects motion intensity. Acceleration $\bm{a}_{i,n}$ is particularly informative for contact-dominant events such as impacts, abrupt stops, or braking, where rapid changes in motion often correlate with stronger and more salient acoustic responses. Its direction indicates whether motion is increasing or decelerating, while its magnitude $\|\bm{a}_{i,n}\|_{2}$ provides a useful cue for event strength.
We use $\|\bm{v}_{i,n}\|_{2}$ and $\|\bm{a}_{i,n}\|_{2}$ as explicit motion-intensity cues that help the model align audio amplitude and temporal structure with the underlying motion. This improves optimization and helps the model distinguish appropriate sound behaviors, such as impact-like transients versus sustained scraping or rolling sounds.
We combine these terms into an 8D kinematic feature for each object and frame:
$\bm{\phi}_{i,n}=\left[\bm{r}_{i,n}^{\top},\bm{v}_{i,n}^{\top},\bm{a}_{i,n}^{\top},\|\bm{v}_{i,n}\|_{2},\|\bm{a}_{i,n}\|_{2}\right]^{\top}\in\mathbb{R}^{8}.$ (4)
This representation explicitly exposes the key kinematic factors relevant to motion-aware sound generation, namely spatial context, interaction mode, and motion intensity. Since these terms may span several orders of magnitude, we apply signed log-compression before encoding. Let
$\bm{d}_{i,n}=\left[\bm{v}_{i,n}^{\top},\bm{a}_{i,n}^{\top},\|\bm{v}_{i,n}\|_{2},\|\bm{a}_{i,n}\|_{2}\right]^{\top},$ (5)
$\tilde{\bm{d}}_{i,n}=\mathrm{sign}(\bm{d}_{i,n})\odot\log_{10}(1+|\bm{d}_{i,n}|).$ (6)
We then encode the normalized kinematic feature as
$\tilde{\bm{\phi}}_{i,n}=\mathcal{E}_{k}\!\left(\mathrm{Norm}\!\left(\left[\bm{r}_{i,n}^{\top},\tilde{\bm{d}}_{i,n}^{\top}\right]^{\top}\right)\right),$ (7)
yielding latent kinematic tokens in $\mathbb{R}^{d}$ that match the embedding dimension of the audio transformer, where $\mathcal{E}_{k}$ denotes a three-layer MLP.
Audio-Kinematic Fusion. The encoded kinematic tokens $\bm{H}_{\text{kin}}=\{\tilde{\bm{\phi}}_{i,n}\}\in\mathbb{R}^{f\times N\times d}$ are injected into the audio branch of the twin-DiT architecture to explicitly modulate audio generation with motion information. Specifically, within each audio transformer block, an auxiliary cross-attention layer is introduced after the self-attention layer, with audio latent states as queries and kinematic tokens as keys and values. We first project $\bm{H}_{\text{kin}}$ to key and value tensors. For clarity, we omit the multi-head notation:
$\bm{K}_{\text{kin}}=\mathrm{Norm}(\bm{H}_{\text{kin}}\bm{W}_{\text{kin}}^{(K)}),\quad\bm{V}_{\text{kin}}=\bm{H}_{\text{kin}}\bm{W}_{\text{kin}}^{(V)}.$ (8)
To preserve temporal alignment, we apply RoPE [42] to the temporal dimension of $\bm{K}_{\text{kin}}$, reshape $\bm{K}_{\text{kin}}$ and $\bm{V}_{\text{kin}}$ to $\mathbb{R}^{(fN)\times d}$, and apply RoPE to audio latents $\bm{H}_{a}\in\mathbb{R}^{L_{a}\times d}$, with temporal indices scaled by $f/L_{a}$, following Ovi. The kinematic-aware update $\Delta\bm{H}_{a}$ is computed as:
$\bm{Q}_{a}$ $=\mathrm{Norm}(\bm{H}_{a}\bm{W}_{a}^{(Q)}),$ (9)
$\Delta\bm{H}_{a}$ $=\mathrm{Softmax}\!\left(\frac{\bm{Q}_{a}\bm{K}_{\text{kin}}^{\top}}{\sqrt{d}}\right)\bm{V}_{\text{kin}}.$ (10)
To prevent kinematic cues from dominating semantic text conditions during early training, we further employ a learnable gating mechanism. The final hidden state $\bm{H}_{a}^{\prime}$ is updated through a residual connection modulated by a sigmoid-gated scalar $\sigma(\gamma)$, where $\gamma$ is a learnable parameter initialized to a negative value:
$\bm{H}_{a}^{\prime}=\bm{H}_{a}+\sigma(\gamma)\cdot\Delta\bm{H}_{a}.$ (11)
This design allows Tora3 to adaptively balance semantic context and motion-aware conditioning, producing audio that is both semantically appropriate and better aligned with the visual motion.
Hybrid Flow Matching
Directly reusing first-frame latents along trajectories is effective for injecting motion cues, but applying the same construction uniformly to all spatial locations can be limiting. We observe that different regions play different roles during denoising. Trajectory-conditioned regions benefit from stronger appearance anchoring to preserve motion guidance, whereas non-trajectory regions benefit from the standard flow to maintain flexible scene refinement. Motivated by this observation, we propose Hybrid Flow Matching, which modifies the probability flow inside the trajectory region $\Omega_{\text{traj}}=\left\{(i,p_{i,n},q_{i,n})\;\middle|\;i\in\{0,\dots,f-1\},\ n\in\{0,\dots,N-1\}\right\}.$ while preserving the standard flow outside it.
Outside $\Omega_{\text{traj}}$, we follow the standard Flow Matching construction on the clean video latent $x_{0}$:
$x_{t}=(1-t)x_{0}+t\epsilon,\qquad v=\epsilon-x_{0},$ (12)
where $t\in[0,1]$ denotes the flow matching timestep, $\epsilon\sim\mathcal{N}(0,I)$ denotes the standard Gaussian endpoint, and $v$ denotes the target velocity field for unconstrained regions.
Inside $\Omega_{\text{traj}}$, we replace the standard noise endpoint with the trajectory-conditioned latent $x^{\mathrm{traj}}$, yielding
$x_{t}=(1-t)x_{0}+tx^{\mathrm{traj}},\qquad v=x^{\mathrm{traj}}-x_{0}.$ (13)
Intuitively, this preserves the standard denoising behavior in unconstrained regions while anchoring trajectory-conditioned regions to the injected motion prior.
To implement this hybrid construction, we use the binary indicator mask $M\in\{0,1\}^{f\times h\times w}$ associated with $\Omega_{\text{traj}}$, defined by $M[i,p,q]=\mathbf{1}[(i,p,q)\in\Omega_{\text{traj}}]$. The resulting latent input and target velocity field are defined as
$x_{t}=(1-M)\odot\big((1-t)x_{0}+t\epsilon\big)+M\odot\big((1-t)x_{0}+tx^{\mathrm{traj}}\big),$ (14)
$v=(1-M)\odot(\epsilon-x_{0})+M\odot(x^{\mathrm{traj}}-x_{0}).$ (15)
Here, the mask $M$ is broadcast along the channel dimension to $\mathbb{R}^{f\times h\times w\times d}$. At inference time, we use the same trajectory-conditioned latent construction to initialize the flow-matching process at $t=1$, with $x_{1}=(1-M)\odot\epsilon+M\odot x^{\mathrm{traj}}$.
This construction ensures that the latent input at each spatial location is consistent with the corresponding velocity target used for supervision. In practice, it helps preserve trajectory-guided motion while reducing local inconsistencies around trajectory-conditioned regions.
During training, we keep the audio objective unchanged from Ovi and only replace its original video loss with a region-balanced objective to better balance the sparse trajectory region and its complement. To ensure smooth boundary handling, we use a soft mask $M_{\text{soft}}$ obtained by Gaussian blurring the binary mask $M$. The modified video loss is defined as
$\mathcal{L}_{\text{video}}=\lambda_{\text{out}}\mathcal{L}_{\text{out}}+\lambda_{\text{traj}}\mathcal{L}_{\text{traj}},$ (16)
where $\lambda_{\text{out}}=\lambda_{\text{traj}}=0.5$. The loss terms are defined as
$\mathcal{L}_{\text{out}}$ $=\frac{\sum\big((1-M_{\text{soft}})\odot(\hat{v}-v)^{2}\big)}{\sum(1-M_{\text{soft}})+\delta},$ (17)
$\mathcal{L}_{\text{traj}}$ $=\frac{\sum\big(M_{\text{soft}}\odot(\hat{v}-v)^{2}\big)}{\sum M_{\text{soft}}+\delta},$ (18)
where $\delta=10^{-8}$ is a small constant for numerical stability. This equal weighting prevents the sparse trajectory region from being overwhelmed by the dominant non-trajectory area, allowing the model to learn localized motion control and global appearance generation in a more balanced manner. Following Ovi, the final training objective combines the video and audio losses with weights $0.85$ and $0.15$, respectively:
$\mathcal{L}_{\text{final}}=0.85\,\mathcal{L}_{\text{video}}+0.15\,\mathcal{L}_{\text{audio}}^{\text{Ovi}},$ (19)
where $\mathcal{L}_{\text{audio}}^{\text{Ovi}}$ denotes the original audio loss used in Ovi.
Table 1: Comparison on video quality, audio quality, text alignment, motion control, and audio-video synchronization, including event-level and intensity-level motion-sound coherence metrics.
Video Quality Audio Quality Text Alignment AV Synchronization Motion Control
Method # Params AS $\uparrow$ FVD $\downarrow$ CE $\uparrow$ CU $\uparrow$ PC $\downarrow$ PQ $\uparrow$ CLAP $\uparrow$ CLIP-T $\uparrow$ FGAS $\uparrow$ ETE $\downarrow$ MAIC $\uparrow$ TE $\downarrow$
LTX-2 [19] 22.16B 4.31 989.6 3.28 6.17 2.43 6.73 0.31 0.29 0.187 0.284 0.41 -
Ovi [31] 11.66B 4.40 887.7 3.30 6.01 1.85 6.44 0.43 0.30 0.156 0.301 0.37 -
MOVA [43] 30.00B 4.63 849.8 3.05 6.31 1.77 6.95 0.46 0.31 0.201 0.236 0.49 -
AVControl [2] 22.32B 4.52 829.6 3.29 6.22 2.18 6.79 0.39 0.30 0.209 0.214 0.55 19.95
Tora3 12.25B 4.61 784.1 3.34 6.43 1.81 7.09 0.44 0.31 0.234 0.181 0.63 12.13
Experiments
Experimental Setup
Datasets. We construct PAV from filtered high-quality subsets of VGGSound [6], ACAV-100M [26], OpenVid1M [33], Pexels [35], and in-house collected data. To improve motion relevance, we use Qwen3-VL [1] to automatically filter clips whose main object exhibits basic motion patterns such as translation, rotation, sliding, parabolic motion, oscillation, scale change, or deformation. For the retained clips, we first apply SAM2 [37] to segment all objects in the first frame, and then run CoTracker3 [22] initialized from the centroid of each object to obtain trajectory annotations, resulting in 460k video clips. These annotations provide scalable kinematic supervision rather than full physical measurements. We further use Qwen3-VL-8B-Instruct and Qwen3-Omni-Captioner [57] to generate paired video and audio descriptions. For quantitative evaluation, we use 50 representative videos covering diverse motion patterns and scene types.
Metrics. We evaluate generation quality from four aspects: video quality, audio quality, text alignment, and audio-video synchronization. For video, we report Fréchet Video Distance (FVD) [45], Aesthetic Score (AS), computed as the average of MANIQA [58] and MUSIQ [23], and Trajectory Error (TE), defined as the L2 distance between the conditioning trajectories and the object tracks estimated from generated videos, following [63]. For audio, we adopt AudioBox-Aesthetics [44], including Content Enjoyment (CE), Content Usefulness (CU), Production Complexity (PC), and Production Quality (PQ). To assess text alignment, we compute CLIP [36] and CLAP [53] scores for text-to-video and text-to-audio consistency, respectively. For audio-video synchronization, we utilize FGAS in PhyAVBench [55], which computes the cosine similarity between video frames and synchronized audio features in the latent space of CAV-MAE Sync [18]. To further assess physical coherence at the motion-sound level, we introduce two additional metrics. Event Timing Error (ETE) measures the time difference between trajectory-derived events and corresponding audio onsets. Motion-Audio Intensity Correlation (MAIC) computes the Pearson correlation between a trajectory-derived motion-intensity signal and the temporally aligned audio energy envelope. We evaluate ETE on clips with discrete events and MAIC on clips with sustained motion and continuous sound evolution. Detailed computation procedures are provided in the appendix.
Implementation Details. The model is initialized from the pretrained Ovi checkpoint and trained for 30k steps on 32 NVIDIA A100 GPUs with a global batch size of 32. We use AdamW with $\beta_{1}=0.9$, $\beta_{2}=0.999$, weight decay 0.01, and learning rate $4\times 10^{-5}$. Training stability is improved with BF16 mixed precision and gradient clipping at 1.0. Kinematic features are normalized using dataset-level statistics computed from 5,000 randomly sampled training examples. We further apply trajectory condition dropout with probability $p=0.05$ to improve robustness. For the mask $M_{\text{soft}}$, the binary trajectory mask is smoothed with a Gaussian kernel of $\sigma=0.5$ to encourage smooth transitions between trajectory-conditioned regions and their surroundings. The sigmoid-gated scalar parameter $\gamma$ is initialized to 10-10.
Main Results
Table 1 compares Tora3 with four competitive baselines, LTX-2 [19], Ovi [31], MOVA [43], and AVControl [2], across video quality, audio quality, text alignment, audio-video synchronization, and motion control. AVControl is a concurrent work on motion-conditioned audio-video generation, and we include its results here for a direct comparison. Overall, Tora3 achieves the best trade-off of generation quality and motion-sound coherence. It delivers the best video realism with the lowest FVD, strong overall audio quality, and the highest FGAS, indicating improved synchronization between generated motion and sound. Tora3 also achieves the lowest ETE and the highest MAIC, showing that it more accurately aligns salient audio events with motion and contact cues, and better matches the evolution of audio intensity to motion strength. At the same time, Tora3 remains highly competitive in text alignment, matching the best CLIP-T score and achieving a strong CLAP score.
Figure 4: Qualitative comparison. Tora3 produces more stable visual motion and more coherent audio in both examples, including motion-consistent loudness variation for the moving motorcycle and better synchronized rolling and collision sounds for the rolling ball. Please refer to the supplementary video for full comparisons.
Compared with AVControl, Tora3 is particularly strong in motion-sound coherence. AVControl mainly uses trajectories as a control signal for the joint backbone, whereas Tora3 uses them as a shared kinematic prior for both modalities. Trajectory-aligned motion injection improves video motion fidelity, and kinematic-audio alignment explicitly links motion states to sound generation, leading to better FGAS, ETE, and MAIC. These gains support our central design: trajectories effectively couple visual motion and acoustic generation under a shared trajectory-grounded structure.
Figure 4 shows the qualitative advantages of Tora3 over competitive baselines. In the motorcycle example, baselines either degrade visual quality or fail to produce motion-consistent loudness variation, whereas Tora3 maintains stable visuals and more coherent audio dynamics. In the rolling-ball example, baselines struggle with rolling, collision, and post-impact motion together with synchronized sound, often producing unrealistic motion, incorrect impact timing, or audio artifacts. Tora3 models these events more faithfully with better synchronized rolling and impact sounds.
Figure 5: Effect of motion speed on generated audio. Tora3 generates audio that varies consistently with motion speed.
Figure 5 shows that Tora3 produces speed-dependent audio variations. When the box is pulled faster, the generated sound is louder and exhibits subtle timbral differences compared with the sound produced under slower motion. This highlights the advantage of using trajectories as a shared kinematic prior, allowing motion changes to be reflected more consistently in both visual dynamics and generated sound.
Ablation Studies
Shared kinematic prior across modalities. Table 2 studies whether trajectory-derived kinematic conditioning should be applied to the video branch, the audio branch, or both. Applying it only to the video branch mainly improves visual generation, yielding the best Aesthetic Score and clearly better video realism. In contrast, applying it only to the audio branch more directly improves audio quality and motion-sound coherence, as reflected by higher PQ, FGAS, and MAIC as well as lower ETE. This asymmetric behavior highlights that the two designs are complementary: trajectory injection improves motion faithfulness in the video branch, while kinematic conditioning improves event timing and intensity alignment in the audio branch. When used together, they produce the strongest overall results in video realism, audio quality, and motion-sound coherence. This confirms that trajectories are most effective as a shared kinematic prior across modalities rather than isolated controls for either branch alone.
Table 2: Effect of using trajectories as a shared kinematic prior for both modalities.
Setting AS $\uparrow$ FVD $\downarrow$ PQ $\uparrow$ FGAS $\uparrow$ ETE $\downarrow$ MAIC $\uparrow$
Neither 4.39 854.7 6.75 0.173 0.296 0.39
Video only 4.51 823.6 6.81 0.198 0.247 0.46
Audio only 4.42 845.2 6.89 0.209 0.221 0.61
Both (full) 4.47 811.8 6.93 0.225 0.193 0.66
Motion representation designs for video-branch injection. Table 3 compares three motion injection strategies for the video branch: a Tora-inspired design using an auxiliary motion encoder and fuser, a WanMove-inspired design that concatenates latent noise with trajectory-guided first-frame propagated features along the channel dimension, and our Tora3 motion representation. Tora3 consistently achieves the best overall performance, yielding the highest visual quality, the strongest audio-video synchronization, the lowest event timing error, and the lowest trajectory error, while requiring no additional parameters. This advantage likely stems from the fact that: unlike prior approaches that transform sparse control signals into separate feature spaces, Tora3 injects motion directly in the native latent space by reusing first-frame features along the prescribed trajectories. This leads to cleaner and more reliable motion control, improving trajectory following in the video branch and providing more stable motion cues to the audio branch through cross-modal interaction.
Kinematic state components. Table 4 studies the effect of progressively enriching the kinematic conditioning used by the audio branch. Starting from no kinematic conditioning, adding position and velocity already improves both synchronization and motion-sound coherence. Incorporating acceleration further strengthens performance, especially in FGAS, ETE, and MAIC. The full second-order kinematic state, which additionally includes velocity and acceleration magnitudes, achieves the best performance across all metrics. This trend is consistent with the role of different kinematic components. Position provides coarse scene context, velocity indicates how objects move over time, and acceleration is particularly informative for abrupt motion changes associated with impact-like events. Adding the magnitudes of velocity and acceleration further exposes motion strength, allowing the model to better calibrate audio intensity and temporal salience. The monotonic improvements therefore support our formulation of trajectory-derived kinematic states as a compact and effective interface between trajectories and audio generation.
Table 3: Ablation on motion representation designs for video-branch injection. Tora-style and WanMove-style denote variants inspired by Tora [63] and WanMove [9], respectively.
Method AS $\uparrow$ FGAS $\uparrow$ ETE $\downarrow$ TE $\downarrow$ # Params
Tora-style [63] 4.49 0.179 0.271 17.06 12.65B (+0.99B)
WanMove-style [9] 4.44 0.184 0.251 13.91 11.66B (+590K)
Ours 4.51 0.198 0.247 13.03 11.66B (+0)
Table 4: Effect of different kinematic state components on audio generation and motion-sound coherence.
Kinematic Signal PQ $\uparrow$ FGAS $\uparrow$ ETE $\downarrow$ MAIC $\uparrow$
None 6.75 0.173 0.296 0.39
$\bm{r}+\bm{v}$ 6.79 0.191 0.254 0.48
$\bm{r}+\bm{v}+\bm{a}$ 6.86 0.202 0.229 0.56
$\bm{r}+\bm{v}+\bm{a}+\|\bm{v}\|_{2}+\|\bm{a}\|_{2}$ 6.89 0.209 0.221 0.61
Table 5: Effect of Hybrid Flow Matching (HFM) on video quality, trajectory following, and overall AV synchronization.
Method AS $\uparrow$ FVD $\downarrow$ TE $\downarrow$ FGAS $\uparrow$
w/o HFM 4.47 811.8 12.94 0.225
w/ HFM 4.61 784.1 12.13 0.234
Hybrid flow matching. Table 5 shows that hybrid flow matching consistently improves both perceptual and distributional video quality, while also improving trajectory faithfulness and overall audio-video synchronization. In particular, it improves the Aesthetic Score from 4.47 to 4.61, reduces FVD from 811.8 to 784.1, and lowers TE from 12.94 to 12.13. These results support our region-aware hybrid construction: stronger appearance anchoring in trajectory-conditioned regions improves motion fidelity, while preserving the standard flow in non-trajectory regions helps maintain local visual coherence.
Conclusion
We presented Tora3, a trajectory-guided framework for audio-video generation that improves physical coherence by using object trajectories as a shared kinematic prior for video and audio synthesis. Tora3 integrates a trajectory-aligned motion representation, a second-order kinematics-based audio alignment module, and a hybrid flow matching objective for robust motion modeling and local consistency. We also introduced PAV, a large-scale audio-video dataset with motion-centric annotations for trajectory-aware generation. Extensive experiments show that Tora3 improves visual realism, audio quality, motion-sound coherence, and audio-video synchronization over strong open-source baselines, while remaining competitive in text alignment. Future work includes extending trajectory-level coherence to richer physical factors, such as material properties, contact dynamics, and 3D acoustic propagation.
[…]
Supplementary Material
Appendix A Overview
This supplementary material provides additional details to complement the main paper. In particular, it includes additional method details, dataset details, evaluation metric details, user study results, limitations, experiment configuration, and notation.
Appendix B Additional Method Details
B.1 Trajectory Processing Pipeline
Given raw object trajectories extracted in image coordinates, we first temporally align them with the latent video frames used by the generator while preserving the first frame. Suppose the original video contains $F$ frames. We retain the trajectory points at the first frame unchanged, and then compute one averaged trajectory point for every subsequent non-overlapping 4-frame interval. Therefore, the number of aligned latent frames is
$f=\left\lfloor\frac{F-1}{4}\right\rfloor+1,$ (20)
which yields
$\mathcal{T}\in\mathbb{R}^{f\times N\times 2},$ (21)
where $N$ is the number of trajectories.
We then map image-space coordinates to latent-space coordinates using the VAE downsampling ratio $s$. For each image-space trajectory point $(j_{i,n},k_{i,n})$, the corresponding latent coordinate $\mathcal{T}^{\prime}_{i,n}=(p_{i,n},q_{i,n})$ is obtained by nearest-neighbor projection onto the latent grid after downsampling by a factor of $s$. In practice, this simple projection is sufficient because the trajectory signal is used as a sparse conditioning prior rather than a hard pixel-level constraint.
B.2 Trajectory Mask and Hybrid Flow Construction
The binary trajectory mask $M\in\{0,1\}^{f\times h\times w}$ marks latent locations visited by trajectory points:
$M[i,p,q]=\mathbf{1}\!\left[(p,q)\in\{(p_{i,n},q_{i,n})\}_{n=1}^{N}\right].$ (22)
Since trajectories are sparse, directly using a hard mask can produce sharp transitions during training. We therefore smooth $M$ with a Gaussian kernel to obtain a soft mask
$M_{\mathrm{soft}}=G_{\sigma}*M,$ (23)
where $G_{\sigma}$ is a Gaussian kernel with standard deviation $\sigma=0.5$. The Gaussian smoothing is applied independently on each frame in the spatial plane. The soft mask is only used in the loss computation; the hybrid flow construction itself still relies on the binary mask. Since we focus on the image-to-video setting, after constructing the input latent $x_{t}$ from $M$, $x_{0}$, $\epsilon$, and $x^{\text{traj}}$, we overwrite the first-frame latent of $x_{t}$ with the given first-frame latent $z$ in both training and inference.
B.3 Boundary Treatment for Kinematic Features
We compute kinematic features using finite differences. For interior frames, velocity and acceleration are estimated using central finite differences, as described in the main paper. Since central differences are not directly applicable at sequence boundaries, we instead adopt second-order accurate forward and backward finite-difference schemes for the first and last frames, respectively:
$\bm{v}_{0,n}$ $\approx\frac{-3\bm{r}_{0,n}+4\bm{r}_{1,n}-\bm{r}_{2,n}}{2\tau},$ (24)
$\bm{v}_{f-1,n}$ $\approx\frac{3\bm{r}_{f-1,n}-4\bm{r}_{f-2,n}+\bm{r}_{f-3,n}}{2\tau},$ (25)
and
$\bm{a}_{0,n}$ $\approx\frac{2\bm{r}_{0,n}-5\bm{r}_{1,n}+4\bm{r}_{2,n}-\bm{r}_{3,n}}{\tau^{2}},$ (26)
$\bm{a}_{f-1,n}$ $\approx\frac{2\bm{r}_{f-1,n}-5\bm{r}_{f-2,n}+4\bm{r}_{f-3,n}-\bm{r}_{f-4,n}}{\tau^{2}}.$ (27)
The dataset filtering step ensures that all retained trajectories are sufficiently long for second-order estimation. For convenience, our implementation uses the findiff library [14].
B.4 Training Objective Details
The final training objective is
$\mathcal{L}_{\mathrm{final}}=0.85\,\mathcal{L}_{\mathrm{video}}+0.15\,\mathcal{L}_{\mathrm{audio}}^{\mathrm{Ovi}}.$ (28)
The video loss is region-balanced:
$\mathcal{L}_{\mathrm{video}}=\lambda_{\mathrm{out}}\mathcal{L}_{\mathrm{out}}+\lambda_{\mathrm{traj}}\mathcal{L}_{\mathrm{traj}},$ (29)
with $\lambda_{\mathrm{out}}=\lambda_{\mathrm{traj}}=0.5$. This balancing is important because the trajectory region occupies only a small fraction of the latent volume. Without reweighting, the training signal from the trajectory region can be dominated by the unconstrained background, weakening motion control. Notably, the loss on the first frame is excluded during training, as this frame is already known.
Appendix C Dataset Details
C.1 Data Sources and Filtering
The PAV collection is assembled from several public audio-video corpora, including VGGSound [6], ACAV-100M [26], OpenVid1M [33], and Pexels [35], together with in-house collected videos. Rather than maximizing category coverage alone, our construction emphasizes clips that are both of sufficient quality and likely to contain visually meaningful motion for trajectory-aware audio-video modeling. Dataset statistics are shown in Table 6.
Table 6: Dataset statistics of PAV.
Category Statistic Value
Scale Total number of clips 460,218
Total duration (hours) 1564.9
Mean clip duration (s) 12.24
Data sources VGGSound 72,054   (15.7%)
ACAV-100M 119,484   (26.0%)
OpenVid1M 106,971   (23.2%)
Pexels 116,678   (25.4%)
In-house collected 45,031   (9.8%)
We first apply a standard quality-control stage to remove clips with obvious defects. Specifically, we discard samples with low visual quality, severe compression artifacts, resolutions that are either too low or excessively high, abrupt shot transitions, prominent subtitles or overlaid graphics, missing or corrupted frames, and audio tracks that are absent, extremely noisy, or otherwise unusable. This stage is intended to filter out common issues that would interfere with reliable visual analysis, motion extraction, or alignment between video and audio.
After this general cleaning stage, we perform a trajectory-oriented dynamic screening step. A vision-language model, Qwen3-VL [1], is used to identify candidate clips whose foreground content exhibits clearly perceivable temporal dynamics. In practice, we retain clips associated with common motion phenomena such as object displacement, descent under gravity, sliding, rolling, periodic motion, spinning, apparent size variation, and non-rigid shape change. This step biases the dataset toward samples where trajectory cues are likely to be informative for downstream audio-video generation, rather than toward largely static or weakly dynamic scenes.
Listing 1: Python function for computing the center of a binary mask.
1import cv2
2import numpy as np
3
4def mask_center(mask):
5 mask01 = (mask > 0).astype(np.uint8)
6 M = cv2.moments(mask01)
7
8 if M["m00"] == 0:
9 return None
10
11 cx = M["m10"] / M["m00"]
12 cy = M["m01"] / M["m00"]
13 return (cx, cy)
C.2 Motion Annotation Extraction
Trajectory annotations are obtained automatically. For each accepted clip, we use SAM2 [37] on the first frame to produce object masks. For each SAM2 mask, we compute the object center as the centroid of the binary foreground region using cv2.moments from OpenCV [4], as shown in LABEL:lst:mask-center. Concretely, given a binary mask, we first convert it to uint8 format and compute its spatial moments. The object center $(c_{x},c_{y})$ is then obtained as
$c_{x}=\frac{m_{10}}{m_{00}},\hskip 28.80008ptc_{y}=\frac{m_{01}}{m_{00}},$ (30)
where $m_{00}$ denotes the mask area and $m_{10},m_{01}$ are the first-order moments. In implementation, we ignore degenerate masks with $m_{00}=0$. We use these mask centers to initialize CoTracker3 [22] and track object motion through time.
We further filter trajectories using simple heuristics to remove unstable cases. In particular, we discard trajectories that are too short in duration, exhibit negligible overall displacement, or are invalid near image boundaries for a substantial portion of the clip. Clips without any remaining valid foreground trajectory after this filtering step are removed. The final dataset contains approximately 460k clips with trajectory annotations that are suitable for large-scale trajectory-aware training.
It is important to note that these annotations are intended as scalable motion proxies rather than precise physical measurements. Since the pipeline relies on automatic segmentation, centroid-based initialization, and generic tracking, errors may still arise in crowded scenes, under strong occlusion, during rapid appearance variation, or for highly deformable objects. In addition, the motion-focused filtering strategy may skew the dataset toward relatively salient and visually recognizable motion patterns.
C.3 Text Annotation
We use Qwen3-VL-8B-Instruct and Qwen3-Omni-Captioner [57] to generate paired text descriptions for video and audio content. The system prompts for generating video and audio descriptions are shown as LABEL:l-video and LABEL:l-audio, respectively. These descriptions are used as text prompts during training and evaluation.
C.4 Test Set Construction
For quantitative evaluation, we build a held-out test set of 50 representative clips that are fully excluded from training. We sample clips to cover diverse motion patterns and scene types, and apply the same quality-control and motion-oriented filtering pipeline as used in the construction of PAV. We further ensure that the selected clips contain clear foreground motion together with either discrete motion-related sound events or sustained motion-dependent audio, so that they are suitable for evaluating trajectory following, event timing, and motion-audio intensity coherence.
Listing 2: The system prompt for video description
1Please provide a detailed, objective, and comprehensive description of this video in one coherent paragraph. Focus only on visible and verifiable visual content, and describe the scene in a way that is suitable for training a text-to-video model.
2
3Your description should include:
4- Main subjects: Identify the primary people, animals, or objects, including their appearance, clothing, pose, materials, colors, and visible textures.
5- Actions and motion: Describe what each subject is doing, including body movement, interactions, motion direction, speed, intensity, and temporal progression.
6- Environment and background: Describe the location, setting, background elements, weather, time of day, lighting conditions, and spatial layout.
7- Camera work and framing: Describe the shot type and composition (e.g. close-up, medium shot, wide shot, overhead view), as well as any camera movement such as pan, tilt, zoom, tracking, handheld motion, or static framing.
8- Visual style and atmosphere: Specify the visual style (e.g. photorealistic, cinematic, documentary, surveillance, animation, 3D render) and the overall visual tone.
9
10Requirements:
11- Write as a single coherent paragraph.
12- Use precise, neutral, literal language.
13- Avoid subjective interpretation, emotional analysis, symbolism, storytelling, or poetic phrasing.
14- Do not infer intentions, backstory, or anything not clearly visible.
15- If the scene changes over time, describe the events in chronological order.
Listing 3: The system prompt for audio description
1# Role
2You are an audio description assistant. Your task is to analyze the audio track of a video and provide a clear, detailed, and faithful description of everything that can be heard.
3
4# Requirements
5- Listen carefully to the entire audio.
6- Describe all audible elements without adding assumptions or visual information.
7- Describe non-speech sounds, such as:
8 - music (genre, mood, tempo)
9 - environmental or ambient sounds
10 - sound effects
11 - actions implied by sound (e.g., footsteps, door opening)
12- If distinct sound sources or segments are present, present them clearly and unambiguously.
13
14# Constraints
15- Do not invent information that is not audible.
16- Do not describe visual content, focus only on what the audio contains.
17- Do not use line breaks in your output.
Appendix D Evaluation Metric Details
In addition to standard video, audio, and text-alignment metrics, we introduce two motion-sound coherence metrics: Event Timing Error (ETE) and Motion-Audio Intensity Correlation (MAIC).
D.1 Event Timing Error (ETE)
Event Timing Error (ETE) measures temporal alignment between visually inferred motion events and salient audio onsets. Because precise event times are not directly available for every generated sample, and exhaustive manual annotation is costly, we use trajectory-based automatic detection to assist manual annotation. We evaluate ETE only on impact-dominant clips, where temporally localized motion events are well defined and more likely to correspond to discrete audio onsets.
Trajectory-derived event candidates. For each generated trajectory, we first smooth the temporal coordinate sequence to reduce high-frequency jitter, and then compute finite-difference accelerations. Candidate visual event times are identified as temporally isolated local extrema in acceleration magnitude $\|\bm{a}_{i,n}\|_{2}$, which capture abrupt kinematic changes. To improve robustness, we retain only extrema whose magnitude exceeds an adaptive threshold defined as the 95th percentile of acceleration magnitudes over all trajectories in the video, and apply temporal non-maximum suppression to remove near-duplicate detections. In impact-dominant motions, such extrema often correlate with contact-like events or sudden momentum changes, making them a useful source of candidate event times.
Manual annotation with automatic assistance. Since trajectory extrema are only a proxy and may contain false positives or miss valid events, the automatically detected candidates are used only to assist annotation rather than define the final event set. Annotators review each clip and annotate the complete set of plausible sound-producing events, including validating or discarding proposed candidates and adding events missed by the automatic procedure. We define such events as visually identifiable, temporally localized motion events that would normally induce salient acoustic responses, such as impacts, contacts, collisions, or abrupt stops. The same event definition and annotation protocol are used for all compared methods. To reduce bias, annotation is performed with model identity hidden from the annotator.
Audio onset extraction. We use the librosa library [32] to detect audio onsets, i.e., time points at which new sound events begin. Specifically, we apply librosa.onset.onset_detect, which computes an onset strength envelope and then performs peak picking over this temporal novelty curve. Intuitively, the onset strength envelope measures abrupt spectral changes in the audio signal, and peaks in this curve often correspond to perceptually salient event boundaries. This provides a standard and robust way to convert a waveform into a sequence of onset timestamps for downstream alignment analysis.
Matching under a tolerance window. Let $\{\theta_{u}^{\mathrm{vis}}\}_{u=1}^{N_{\mathrm{vis}}}$ denote the filtered visual event times, and let $\{\theta_{l}^{\mathrm{audio}}\}_{l=1}^{N_{\mathrm{audio}}}$ denote the detected audio onset times. We seek an optimal one-to-one matching between visual and audio events under a temporal tolerance threshold $\Delta_{\mathrm{tol}}$. We use $\Delta_{\mathrm{tol}}=0.5$ s as a conservative matching window. This value is large enough to tolerate approximation errors in automatically derived visual events and audio onset detection, while remaining informative in practice. Let $\Pi$ denote any one-to-one matching, where each visual event and each audio event can appear in at most one matched pair. We regard $\Pi$ as valid only if:
$\left|\theta_{u}^{\mathrm{vis}}-\theta_{l}^{\mathrm{audio}}\right|\leq\Delta_{\mathrm{tol}},\hskip 28.80008pt\forall(u,l)\in\Pi.$ (31)
Error computation. For a given valid matching $\Pi$, define the number of unmatched events as
$N_{\mathrm{unm}}=(N_{\mathrm{vis}}-|\Pi|)+(N_{\mathrm{audio}}-|\Pi|).$ (32)
Each matched pair contributes its absolute temporal deviation, while each unmatched event incurs a penalty of $\Delta_{\mathrm{tol}}$. The matching-specific error is defined as
$\mathrm{ETE}(\Pi)=\frac{\sum_{(u,l)\in\Pi}\left|\theta_{u}^{\mathrm{vis}}-\theta_{l}^{\mathrm{audio}}\right|+\Delta_{\mathrm{tol}}\,N_{\mathrm{unm}}}{|\Pi|+N_{\mathrm{unm}}}.$ (33)
The final event timing error is then defined as the minimum over all valid one-to-one matchings:
$\mathrm{ETE}=\min_{\Pi}\mathrm{ETE}(\Pi).$ (34)
This formulation jointly balances temporal alignment quality and penalties for unmatched events. Lower ETE indicates better temporal alignment between visually inferred impact events and generated sound.
As a boundary convention, when both the filtered visual event set and the detected audio onset set are empty, we set
$\mathrm{ETE}=\Delta_{\mathrm{tol}}.$ (35)
Although this case yields no explicit timing mismatch, we treat it as a failure case rather than perfect alignment, since ETE is evaluated only on impact-dominant clips where at least one salient event is generally expected. This prevents degenerate no-event generations from receiving an artificially optimal score.
D.2 Motion-Audio Intensity Correlation (MAIC)
Motion-Audio Intensity Correlation (MAIC) evaluates whether the temporal evolution of audio intensity is consistent with the temporal evolution of visible motion strength.
Trajectory-derived motion-strength signal. For each object trajectory, we compute a scalar motion-strength signal at frame $i$ by combining the velocity magnitude and the acceleration magnitude:
$\mu_{i,n}=\|\bm{v}_{i,n}\|_{2}+\|\bm{a}_{i,n}\|_{2}.$ (36)
This simple combination captures both sustained motion magnitude and temporal changes in motion. When multiple trajectories are present, we directly average the motion-strength signals across objects to obtain a clip-level motion-strength sequence, without introducing additional object-specific weighting:
$\mu_{i}=\frac{1}{N}\sum_{n=1}^{N}\mu_{i,n},\hskip 28.80008pti=0,\dots,f-1.$ (37)
Audio-intensity signal. Given a generated waveform sampled at 16 kHz, we compute a frame-level log-energy signal aligned with the motion-strength sequence $\{\mu_{i}\}_{i=0}^{f-1}$. Let $N_{\mathrm{wav}}$ denote the total number of waveform samples in the generated clip, and let $y[i_{\mathrm{wav}}]$ denote the waveform sample at discrete sample index $i_{\mathrm{wav}}$. We divide the waveform into $f$ non-overlapping temporal segments aligned with the video frames. Specifically, for each frame $i=0,\dots,f-1$, we define
$\mathcal{I}_{i}=\left[\frac{iN_{\mathrm{wav}}}{f},\,\frac{(i+1)N_{\mathrm{wav}}}{f}\right)\cap\mathbb{Z},\hskip 28.80008pti=0,\dots,f-1.$ (38)
We compute the mean squared energy of the $i$-th segment as
$e_{i}=\frac{1}{|\mathcal{I}_{i}|}\sum_{i_{\mathrm{wav}}\in\mathcal{I}_{i}}y[i_{\mathrm{wav}}]^{2}.$ (39)
To reduce dynamic range and improve numerical stability, we then apply logarithmic compression:
$\alpha_{i}=\log(1+e_{i}),\hskip 28.80008pti=0,\dots,f-1.$ (40)
This yields a frame-aligned audio-intensity sequence $\{\alpha_{i}\}_{i=0}^{f-1}$.
Correlation computation. MAIC is defined as the Pearson correlation coefficient between the clip-level motion-strength sequence $\{\mu_{i}\}_{i=0}^{f-1}$ and the audio-intensity sequence $\{\alpha_{i}\}_{i=0}^{f-1}$:
$\mathrm{MAIC}=\frac{\sum_{i=0}^{f-1}(\mu_{i}-\bar{\mu})(\alpha_{i}-\bar{\alpha})}{\sqrt{\sum_{i=0}^{f-1}(\mu_{i}-\bar{\mu})^{2}}\sqrt{\sum_{i=0}^{f-1}(\alpha_{i}-\bar{\alpha})^{2}}},$ (41)
where
$\bar{\mu}=\frac{1}{f}\sum_{i=0}^{f-1}\mu_{i},\hskip 28.80008pt\bar{\alpha}=\frac{1}{f}\sum_{i=0}^{f-1}\alpha_{i}$ (42)
are the temporal means of the motion-strength and audio-intensity signals, respectively. Higher MAIC indicates stronger agreement between motion variation and audio-energy variation over time. If either sequence has zero temporal variance, we set MAIC to 0.
Usage protocol. MAIC is most informative for clips with sustained motion and continuously varying sound, such as driving, dragging, sliding, or rolling. In contrast to ETE, which focuses on temporally localized impact-like events, MAIC captures whether stronger visible motion is accompanied by correspondingly stronger generated sound over the course of a clip.
Table 7: User study results (%) in pairwise comparison against baselines. Each entry is reported as Win / Tie / Lose for Tora3. Higher win rate indicates stronger human preference for Tora3. Motion Alignment is only applicable to methods that support explicit trajectory control.
Tora3 compared to VQ EAT MSIA MA OP
LTX-2 59.2 / 15.9 / 24.9 60.1 / 16.7 / 23.2 63.9 / 21.2 / 14.9 - 59.1 / 16.8 / 24.1
Ovi 54.4 / 16.8 / 28.8 65.9 / 18.2 / 15.9 66.1 / 21.8 / 12.1 - 63.9 / 18.1 / 18.1
MOVA 48.8 / 20.7 / 30.5 51.1 / 21.9 / 27.0 50.8 / 26.3 / 22.9 - 52.3 / 22.8 / 24.9
AVControl 52.6 / 19.8 / 27.6 48.4 / 22.7 / 29.0 44.1 / 28.9 / 27.0 45.1 / 23.9 / 31.0 50.2 / 23.1 / 26.7
Figure 6: Pairwise user study results of Tora3 against each baseline, reported as win–tie–lose rates (%). Motion Alignment is only applicable to methods that support explicit trajectory control.
Appendix E User Study
Automatic metrics quantify different aspects of generation quality, but they do not fully capture how users perceive motion-conditioned audio-video generation. We therefore conduct a human preference study to examine whether the advantages of Tora3 are also meaningful at the perceptual level. Instead of asking participants to rate abstract metric dimensions, we ask them to compare models according to five concrete criteria: Video Quality, Event-level Audio Timing, Motion-Sound Intensity Alignment, Motion Alignment, and Overall Preference.
Setup. We compare Tora3 against four baselines: LTX-2, Ovi, MOVA, and AVControl. We recruit 10 participants for the study, and each participant evaluates 100 comparison pairs, resulting in 1000 pairwise comparisons in total. The evaluated pairs are uniformly distributed across model pairs, so that each participant reviews a balanced number of comparisons for each baseline. For each pairwise comparison, participants are shown two generated clips under the same text prompt and trajectory condition (when supported), with model identities hidden and left-right order randomized. They are then asked to compare the two anonymized videos for each of the following questions and assign a label of win, tie, or lose:
  • Video Quality (VQ): Which video looks better in terms of naturalness, temporal stability, and visual coherence?
  • Event-level Audio Timing (EAT): Which video better aligns salient sound events (e.g., impact, contact, abrupt motion change) with the visible motion?
  • Motion-Sound Intensity Alignment (MSIA): Which video better matches the temporal variation of sound intensity to the visible motion intensity over time?
  • Motion Alignment (MA): Which video better follows the prescribed object trajectory and intended motion pattern?
  • Overall Preference (OP): Which video is preferred overall when considering both video and audio together?
These perceptual criteria are designed to correspond to the major dimensions in our quantitative evaluation. This design allows us to directly assess whether the quantitative improvements of Tora3 translate into a better user experience.
Analysis. The user study results in Table 7 and Figure 6 show that the advantages of Tora3 are consistently reflected in human perception. Across most pairwise comparisons and evaluation dimensions, participants tend to prefer Tora3 over the competing methods, indicating that its improvements are not only measurable by automatic metrics but also perceptually meaningful.
A particularly clear trend appears on Event-level Audio Timing (EAT) and Motion-Sound Intensity Alignment (MSIA). Participants generally judge Tora3 to better place salient sound events at visually appropriate moments and to better match the evolution of sound intensity with the evolution of visible motion strength. This suggests that the proposed trajectory-aware audio-video generation framework improves both localized synchronization, such as impact timing, and more continuous correspondence between motion dynamics and audio energy. These perceptual findings are well aligned with the motivation behind our proposed coherence metrics.
Tora3 is also generally preferred in terms of Video Quality (VQ). Although the margin is not as pronounced as in the motion-sound criteria, the overall preference trend indicates that introducing trajectory-conditioned control and audio-aware modeling does not harm visual realism or temporal consistency. Instead, users often perceive the resulting videos as more natural and coherent. This is important because it suggests that the gains in controllability and cross-modal alignment are achieved without sacrificing visual quality.
For Motion Alignment (MA), which is only applicable in comparison with AVControl, participants show a moderate preference for Tora3 on motion alignment and overall preference, while the margin is smaller than against non-trajectory-controlled baselines. This indicates that Tora3 more faithfully follows the prescribed trajectory and intended motion pattern in a perceptually meaningful way.
The trend in Overall Preference (OP) further confirms the above observations. When users are asked to consider the generated video and audio jointly, Tora3 is generally preferred over all baselines. This suggests that the benefits of improved event timing, better motion-sound intensity correspondence, and stronger motion controllability combine into a more favorable overall viewing experience. In other words, the proposed method does not merely improve isolated aspects of generation, but leads to a more coherent multimodal result as perceived by human evaluators.
Overall, the human study supports the main claim of our work: trajectory-aware generation improves not only motion controllability, but also the perceptual coherence between visible motion and generated sound. The fact that these advantages are consistently recognized by users provides complementary evidence to the objective evaluation and highlights the practical value of the proposed framework.
Appendix F Limitations
Although Tora3 improves physical coherence, it still has several limitations.
First, trajectories provide only a lightweight kinematic prior. They do not encode material properties, contact stiffness, object mass, surface roughness, or 3D geometry. As a result, Tora3 can improve temporal alignment and intensity evolution but cannot fully determine the exact sound texture of all interactions.
Second, our current formulation uses monaural audio generation. This limits the modeling of spatial acoustics such as direction-dependent attenuation, reverberation, and binaural cues.
Appendix G Experiment Configuration
The main experimental settings are summarized below:
  • Base model: pretrained Ovi “720x720_5s” checkpoint.
  • Training steps: 30,00030{,}000.
  • Hardware: $32$ NVIDIA A100 GPUs.
  • Global batch size: $32$.
  • Video resolution: $960\times 544$.
  • Audio sample rate: $16$ kHz.
  • Number of frames: $121$.
  • Target FPS: $24$.
  • Optimizer: AdamW.
  • Learning rate: $4\times 10^{-5}$.
  • Learning rate schedule: constant.
  • Weight decay: $0.01$.
  • Mixed precision: BF16.
  • Gradient clipping: $1.0$.
  • Trajectory dropout probability: $0.05$.
  • Soft-mask Gaussian smoothing ($\sigma$): $0.5$.
  • Loss weights: video $0.85$, audio $0.15$.
  • Region-balanced video loss weights: $\lambda_{\mathrm{out}}=\lambda_{\mathrm{traj}}=0.5$.
  • Numerical stability constant ($\delta$): $10^{-8}$.
  • Inference sampling steps: $50$.
Appendix H Notations
$x_{0}$: clean video latent sequence
$z$: first-frame latent
$\widetilde{\mathcal{T}}$: complete object trajectory in pixel space
$\mathcal{T}$: trajectory tensor aligned with the sampled video frames
$\mathcal{T}^{\prime}$: trajectory mapped to latent grid coordinates
$x^{\mathrm{traj}}$: trajectory-conditioned latent constructed under trajectory guidance
$x_{t}$: intermediate latent at time $t$ in flow matching
$x_{1}$: initialization latent at $t=1$ during inference
$\epsilon$: standard Gaussian noise endpoint
$v$: target velocity field in flow matching
$\hat{v}$: model-predicted velocity field
$h$: spatial height of the latent
$w$: spatial width of the latent
$d$: channel or embedding dimension of the latent / token
$f$: number of latent video frames
$F$: number of original video frames
NN: number of trajectories
$i$: frame index
$n$: index of the $n$-th trajectory
$p_{i,n}$: row coordinate of the $n$-th trajectory point in frame $i$ on the latent grid
$q_{i,n}$: column coordinate of the $n$-th trajectory point in frame $i$ on the latent grid
$j_{i,n}$: vertical coordinate in image space
$k_{i,n}$: horizontal coordinate in image space
$s$: VAE downsampling factor
$\bm{r}_{i,n}$: normalized position vector of the $n$-th object in frame $i$
$\bm{v}_{i,n}$: velocity vector of the $n$-th object in frame $i$
$\bm{a}_{i,n}$: acceleration vector of the $n$-th object in frame $i$
$\|\bm{v}_{i,n}\|_{2}$: velocity magnitude
$\|\bm{a}_{i,n}\|_{2}$: acceleration magnitude
$\bm{\phi}_{i,n}$: 8-dimensional kinematic feature: position, velocity, acceleration, and their magnitudes
$\bm{d}_{i,n}$: concatenated dynamic-term vector excluding position
$\tilde{\bm{d}}_{i,n}$: dynamic term after signed log compression
$\tilde{\bm{\phi}}_{i,n}$: normalized and encoded kinematic token
$\mathcal{E}_{k}$: kinematic encoder
$\tau$: time interval between adjacent video frames, $\tau=1/\mathrm{fps}$
$\bm{H}_{\text{kin}}$: kinematic token sequence
$\bm{H}_{a}$: audio token hidden states
$\bm{H}_{a}^{\prime}$: audio hidden states fused with kinematic information
$\Delta\bm{H}_{a}$: incremental update produced by kinematic cross-attention
$\bm{Q}_{a}$: query in the audio-branch cross-attention
$\bm{K}_{\text{kin}}$: key in the kinematic cross-attention
$\bm{V}_{\text{kin}}$: value in the kinematic cross-attention
$L_{a}$: length of the audio tokens
$\bm{W}_{a}^{(Q)}$: projection matrix for generating audio queries
$\bm{W}_{\text{kin}}^{(K)}$: projection matrix for generating kinematic keys
$\bm{W}_{\text{kin}}^{(V)}$: projection matrix for generating kinematic values
$\gamma$: learnable gating parameter
$\sigma(\cdot)$: sigmoid function
$\Omega_{\text{traj}}$: trajectory-conditioned region, corresponding to the set of all latent positions traversed by trajectories
$M$: binary trajectory mask
$M_{\text{soft}}$: soft mask after Gaussian smoothing
$\odot$: Hadamard element-wise multiplication
$\mathbf{1}[\cdot]$: indicator function
$\sigma$: standard deviation of the Gaussian smoothing kernel
$G_{\sigma}$: Gaussian kernel with standard deviation $\sigma$
*: convolution operation
$\mathcal{L}_{\text{video}}$: training loss for the video branch
$\mathcal{L}_{\text{audio}}^{\text{Ovi}}$: original audio loss of Ovi
$\mathcal{L}_{\text{final}}$: final joint training objective
$\mathcal{L}_{\text{out}}$: regression loss for outside of trajectory regions
$\mathcal{L}_{\text{traj}}$: regression loss for trajectory regions
$\lambda_{\text{out}}$: loss weight for outside of trajectory regions
$\lambda_{\text{traj}}$: loss weight for trajectory regions
$\delta$: numerical stability constant
$t$: time step in flow matching, $t\in[0,1]$
$\beta_{1}$: first-moment hyperparameter of AdamW
$\beta_{2}$: second-moment hyperparameter of AdamW
$p$: trajectory condition dropout probability
$(j_{i,n},k_{i,n})$: 2D coordinate of the $n$-th trajectory point in frame $i$ in image space
$\theta_{u}^{\mathrm{vis}}$: timestamp of the $u$-th filtered visual event
$\theta_{l}^{\mathrm{audio}}$: onset time of the $l$-th audio event
$u$: visual event index
$l$: audio event index
$N_{\mathrm{vis}}$: number of visual events
$N_{\mathrm{audio}}$: number of audio events (onsets)
$\Delta_{\mathrm{tol}}$: temporal tolerance window for event matching in ETE
$\Pi$: one-to-one matching set between visual and audio events that satisfies the tolerance-window constraint
$N_{\mathrm{unm}}$: total number of unmatched events
$\mathrm{ETE}(\Pi)$: event temporal error computed under a given matching $\Pi$
$c_{x}$: horizontal coordinate of the centroid of the binary mask
$c_{y}$: vertical coordinate of the centroid of the binary mask
$m_{00}$: zeroth-order moment of the binary mask, corresponding to the foreground area
$m_{10}$: first-order spatial moment of the binary mask (in the $x$ direction)
$m_{01}$: first-order spatial moment of the binary mask (in the $y$ direction)
$\mu_{i,n}$: motion-strength scalar of the $n$-th trajectory in frame $i$
$\mu_{i}$: clip-level motion-strength at frame $i$
$N_{\mathrm{wav}}$: total number of waveform samples in the generated audio clip
$y[i_{\mathrm{wav}}]$: waveform sample at discrete audio sample index $i_{\mathrm{wav}}$
$\mathcal{I}_{i}$: set of waveform sample indices assigned to frame $i$
$e_{i}$: mean squared audio energy of the waveform segment aligned with frame $i$
$\alpha_{i}$: log-compressed audio-intensity at frame $i$
$\bar{\mu}$: temporal mean of the clip-level motion-strength sequence
$\bar{\alpha}$: temporal mean of the audio-intensity sequence