### 16 Apr 2026 v2

# Tora3: Trajectory-Guided Audio-Video Generation with Physical Coherence

Abstract
Audio-video (AV) generation has recently made strong progress in perceptual quality and multimodal coherence, yet generating content with plausible motion-sound relations remains challenging. Existing methods often produce object motions that are visually unstable and sounds that are only loosely aligned with salient motion or contact events, largely because they lack an explicit motion-aware structure shared by video and audio generation. We present Tora3, a trajectory-guided AV generation framework that improves physical coherence by using object trajectories as a shared kinematic prior. Rather than treating trajectories as a video-only control signal, Tora3 uses them to jointly guide visual motion and acoustic events. Specifically, we design a trajectory-aligned motion representation for video, a kinematic-audio alignment module driven by trajectory-derived second-order kinematic states, and a hybrid flow matching scheme that preserves trajectory fidelity in trajectory-conditioned regions while maintaining local coherence elsewhere. We further curate PAV, a large-scale AV dataset emphasizing motion-relevant patterns with automatically extracted motion annotations. Extensive experiments show that Tora3 improves motion realism, motion-sound synchronization, and overall AV generation quality over strong open-source baselines.
Figure 1: Examples generated by Tora3. Tora3 generates audio-video content that follows prescribed trajectories with improved physical coherence, producing more plausible motion and better motion-sound alignment. Long trajectories are segmented for clarity. Please refer to the supplementary video for more examples.
Introduction
Recent advances in generative modeling have substantially improved video synthesis across multiple modalities, including text-to-video (T2V) [43, 3, 23, 59], audio-to-video (A2V) [7, 37, 46], and video-to-audio (V2A) [6, 14, 13]. More recently, unified text-to-audio-video (T2AV) models [17, 29, 31, 9, 11] have emerged, aiming to jointly synthesize visual and acoustic content from a single prompt. Commercial systems such as Veo3 [11], Sora2 [31], and Wan2.5 [9] suggest that high-fidelity multimodal generation is becoming increasingly practical.
Despite this progress, current T2AV generation is still driven primarily by semantic alignment rather than explicit motion-sound modeling. As a result, even strong open-source models [17, 29, 44] often produce visually implausible motion or audio that is only weakly synchronized with salient physical events [19]. In the visual domain, generated objects may follow inconsistent trajectories, exhibit implausible falling patterns, or undergo abrupt accelerations that break temporal continuity [12, 4]. In the acoustic domain, generated sound is often detached from the underlying motion. Impact sounds may miss the moment of contact, sound intensity may fail to reflect interaction strength, and continuous sounds such as sliding or driving often do not evolve with changes in motion. These failures are especially noticeable in interaction-heavy scenes and reduce realism. Existing T2AV methods typically align audio and video at a high semantic level, but lack an explicit intermediate representation that constrains both modalities under a shared motion-aware structure.
In this work, we focus on a practical notion of physical coherence in audio-video generation: whether generated object motion follows plausible trajectory-level dynamics, whether salient sound events are temporally aligned with motion or contact, and whether audio intensity evolves coherently with motion strength. We argue that object trajectories provide an effective and practical interface for introducing such structure into audio-video generation. Trajectories compactly describe where objects are and how they move over time, while exposing kinematic cues such as event timing, motion intensity, and temporal evolution. From this perspective, trajectories offer a natural shared representation for both modalities. They constrain plausible motion in video and indicate when sounds should occur and how their amplitude and temporal structure should evolve. Although trajectories do not fully specify real-world physics, they provide a lightweight and interpretable prior for improving motion-sound coherence in multimodal generation. Figure 2 shows that trajectory guidance better aligns visual motion with corresponding audio events, improving cross-modal synchronization.
[…]
Motivated by these observations, we present Tora3, a trajectory-guided audio-video generation framework that improves physical coherence through shared kinematic conditioning (see Figure 1 for examples). Rather than using trajectories as an auxiliary control signal, Tora3 treats them as a shared kinematic prior to jointly guide video motion and audio generation. Tora3 has three key components. First, for the video branch, we propose an efficient trajectory-aligned motion representation that reuses first-frame latents along sparse trajectories to inject motion cues directly into the latent space without an explicit motion encoder. This design avoids unnecessary distribution shift and preserves consistent motion conditioning throughout the AV pipeline. Second, for the audio branch, we design a kinematic-audio alignment module. Inspired by second-order motion representations [28, 57], we derive compact kinematic descriptors, including position, velocity, acceleration, and their magnitudes, from trajectories and inject them into the audio diffusion model via cross-attention. These trajectory-derived kinematic states provide direct cues for event timing and motion intensity, enabling sound generation whose loudness and temporal evolution better follow the underlying motion. Third, we introduce a hybrid flow matching mechanism with separate probability flows for trajectory and background regions. This mechanism preserves faithful trajectory-guided motion during early denoising while maintaining local appearance coherence during later refinement. To support training and evaluation, we further curate PAV, a large-scale audio-video dataset containing 460k video clips that emphasize motion-relevant patterns and provide automatically extracted motion annotations for trajectory-aware and kinematics-aware AV generation.
Our main contributions are as follows:
  • We propose Tora3, a trajectory-guided audio-video generation framework that uses object trajectories as a shared kinematic prior to jointly guide visual motion and acoustic events, improving physical coherence in generated AV content.
  • We develop a unified trajectory-grounded generation framework with three key components: (1) a trajectory-aligned motion representation for direct video latent conditioning, (2) a kinematic-audio alignment module based on trajectory-derived position, velocity, and acceleration cues, and (3) a hybrid flow matching scheme that preserves motion fidelity while maintaining local consistency.
  • We build PAV, a 460k-clip motion-centric AV dataset with automatic motion annotations, and show through extensive experiments that Tora3 consistently outperforms competitive baselines in video quality, trajectory faithfulness, and motion-sound coherence.
Related work
Trajectory-Controlled Video Generation
Trajectory-controlled video generation has attracted increasing attention for enabling fine-grained manipulation of object motion and scene dynamics. Prior works [26, 38, 56, 48, 25, 60, 8] inject motion cues such as optical flow [56, 38], point tracks [48, 60, 61], or temporally varying bounding boxes [10, 25] into diffusion models via dedicated encoders, adapters, or attention-based fusion modules. To better represent complex motions, some methods [51, 45] further explore 3D trajectories or depth-augmented keypoint maps, while others [15, 8] directly use pretrained base models without architectural modifications to improve scalability. Overall, these approaches substantially improve visual motion controllability and realism under trajectory guidance. However, existing trajectory-controlled methods mainly target video-only generation and do not address joint audio-video synthesis. In contrast, Tora3 extends trajectory guidance to unified audio-video generation by using trajectories as a shared kinematic prior for both motion control and audio-motion alignment.
Audio-Video Generation
Most prior unified audio-video generation methods [27, 36, 47, 53, 58] synthesize ambient sounds using dual-branch architectures, in which audio and video streams are modeled separately and fused through cross-attention or feature concatenation. Large-scale systems such as Veo3 show strong text-conditioned joint generation quality, and recent open-source models such as UniVerse-1 [44], Ovi [29], and LTX-2 [17] largely follow the same dual-backbone fusion paradigm. JoVA [18] further enables direct cross-modal interaction for human-centric content. Nevertheless, these methods often struggle to generate dynamic sequences in which audio remains well aligned with object motion and interaction timing. In contrast, Tora3 introduces trajectory-derived kinematic conditioning into joint AV generation. Specifically, we use object trajectories as a shared kinematic prior, combining trajectory-guided motion control in the video branch with kinematic-audio alignment in the audio branch. This design improves motion-sound coherence by aligning sound events and intensity changes with trajectory-level motion dynamics.
Figure 3: The overall architecture of Tora3. Object trajectories provide a shared kinematic prior for both video and audio generation. The video branch uses trajectory-aligned motion injection, while the audio branch incorporates trajectory-derived kinematic features to better align sound with motion. Hybrid flow matching further improves trajectory-guided generation by balancing motion fidelity and local coherence.
Methods
Overview
Tora3 is an audio-video generation framework built on a twin diffusion transformer architecture with separate video and audio DiT backbones, following Ovi [29]. It uses object trajectories as a shared kinematic prior across both modalities and consists of three components: trajectory-aligned motion representation for video, kinematic-audio alignment module driven by second-order kinematic states for audio, and hybrid flow matching for preserving motion fidelity and local coherence. The overall framework is illustrated in Figure 3.
Trajectory-aligned Motion Representation
Following Tora [60] and Tora2 [61], we use object trajectories to represent both local motion cues and longer-range temporal evolution. A common strategy in controllable video generation is to first encode trajectories with an auxiliary motion encoder and then inject the resulting features into the generative backbone. However, directly adopting such a design in a unified audio-video framework is suboptimal. Since both audio and video branches are large DiT backbones, introducing an additional motion encoder would further entangle optimization and make it harder to preserve a stable motion representation across modalities. In addition, extra encoding layers may attenuate sparse trajectory signals before they reach the backbone, which can reduce motion fidelity.
To avoid these issues, Tora3 directly maps trajectories into the latent space of the video backbone without introducing a dedicated motion encoder. We follow a standard image-to-video pipeline. Let $z\in\mathbb{R}^{h\times w\times d}$ denote the latent of the first frame, obtained by a pretrained VAE encoder [22], where $h$ and $w$ are the latent spatial dimensions and $d$ is the channel dimension. Let $x_{0}\in\mathbb{R}^{f\times h\times w\times d}$ denote the clean video latent sequence, where $f$ is the number of latent frames and the first frame corresponds to the conditioning image. Let $\widetilde{\mathcal{T}}$ denote the full object trajectories in pixel space. We apply temporal average pooling to $\widetilde{\mathcal{T}}$ to match the sampled video frames, obtaining $\mathcal{T}\in\mathbb{R}^{f\times N\times 2}$. We further map $\mathcal{T}$ to latent coordinates by scaling with the VAE downsampling factor and applying nearest-neighbor rounding, yielding $\mathcal{T}^{\prime}$, where each $\mathcal{T}^{\prime}_{i,n}=(p_{i,n},q_{i,n})\in\{0,\dots,h-1\}\times\{0,\dots,w-1\}$. Here, $p_{i,n}$ and $q_{i,n}$ denote the row and column index, respectively, of object $n$ at frame $i$ on the latent grid. We construct a trajectory-conditioned latent $x^{\mathrm{traj}}\in\mathbb{R}^{f\times h\times w\times d}$ by propagating first-frame object features along the trajectories. Specifically, for each object trajectory, we replace the latent feature at the trajectory-aligned location in frame $i$ with the corresponding first-frame feature:
$x^{\mathrm{traj}}[i,p_{i,n},q_{i,n},:]=z[p_{0,n},q_{0,n},:],\quad i\in\{1,\dots,f-1\}.$ (1)
If multiple trajectories map to the same latent location, we randomly select one. All non-trajectory locations in $x^{\mathrm{traj}}$ remain zero, except in the first frame, where they are overwritten by $z$. This design is motivated by two empirical observations: First, VAE latents are spatially smooth, so nearby spatial locations tend to share similar local appearance statistics [35]. Second, relocating latent features to nearby trajectory-aligned positions often preserves local object cues [49]. Together, these observations make first-frame latent propagation a direct and effective way to inject motion cues into the native latent space. This strategy provides an explicit trajectory-aligned conditioning signal while avoiding the additional feature transformation introduced by a separate motion encoder.
Kinematic-Audio Alignment
Semantic conditioning alone is insufficient for realistic motion-dependent audio synchronization. In monaural audio generation, motion-sound alignment depends heavily on temporal structure, spectral texture, and amplitude evolution, all of which are often correlated with object motion. Without explicit kinematic cues, the model must infer these relations implicitly from visual features, often leading to inaccurate timing and mismatched intensity.
Kinematic Feature Representation. To introduce explicit motion-aware structure into audio generation, we condition the audio branch on compact kinematic features derived from object trajectories. These features form a second-order kinematic state, since an object’s motion can be naturally described by its position, velocity, and acceleration. Together, they provide informative cues about when interactions occur, how strong they are, and how acoustic energy may evolve over time.
Given temporal-downsampled trajectories $\mathcal{T}\in\mathbb{R}^{f\times N\times 2}$, we normalize the 2D coordinates by the longer side of the image and denote the normalized position of object $n$ at frame $i$ as $\bm{r}_{i,n}\in[0,1]^{2}$. These normalized positions provide a coarse spatial anchor and useful scene context, such as whether an object is close to the ground or suspended in the air, which may help infer the underlying interaction type and associated sound.
We estimate velocity and acceleration using second-order central differences for interior frames, with forward and backward differences at the boundaries. Let $\tau=1/\text{fps}$ denote the time interval between consecutive frames. The velocity and acceleration at frame $i$ for trajectory point $n$ are approximated as
$\bm{v}_{i,n}\approx\frac{\bm{r}_{i+1,n}-\bm{r}_{i-1,n}}{2\tau},$ (2)
$\bm{a}_{i,n}\approx\frac{\bm{r}_{i+1,n}-2\bm{r}_{i,n}+\bm{r}_{i-1,n}}{\tau^{2}},$ (3)
valid for interior frames $i\in\{1,\dots,f-2\}$. This provides stable, second-order accurate estimates of trajectory-level motion. We explicitly retain both vector components and magnitudes to disentangle direction from strength.
The velocity vector characterizes the mode of interaction, for example, predominantly vertical motion for falling and impact, or horizontal motion for sliding and rolling, which can be related to different sound patterns. Its magnitude $\|\bm{v}_{i,n}\|_{2}$ reflects motion intensity. Acceleration $\bm{a}_{i,n}$ is particularly informative for contact-dominant events such as impacts, abrupt stops, or braking, where rapid changes in motion often correlate with stronger and more salient acoustic responses. Its direction indicates whether motion is increasing or decelerating, while its magnitude $\|\bm{a}_{i,n}\|_{2}$ provides a useful cue for event strength.
We use $\|\bm{v}_{i,n}\|_{2}$ and $\|\bm{a}_{i,n}\|_{2}$ as explicit motion-intensity cues that help the model align audio amplitude and temporal structure with the underlying motion. This improves optimization and helps the model distinguish appropriate sound behaviors, such as impact-like transients versus sustained scraping or rolling sounds.
We combine these terms into an 8D kinematic feature for each object and frame:
$\bm{\phi}_{i,n}=\left[\bm{r}_{i,n}^{\top},\bm{v}_{i,n}^{\top},\bm{a}_{i,n}^{\top},\|\bm{v}_{i,n}\|_{2},\|\bm{a}_{i,n}\|_{2}\right]^{\top}\in\mathbb{R}^{8}.$ (4)
This representation explicitly exposes the key kinematic factors relevant to motion-aware sound generation, namely spatial context, interaction mode, and motion intensity. Since these terms may span several orders of magnitude, we apply signed log-compression before encoding. Let
$\bm{d}_{i,n}=\left[\bm{v}_{i,n}^{\top},\bm{a}_{i,n}^{\top},\|\bm{v}_{i,n}\|_{2},\|\bm{a}_{i,n}\|_{2}\right]^{\top},$ (5)
$\tilde{\bm{d}}_{i,n}=\mathrm{sign}(\bm{d}_{i,n})\odot\log_{10}(1+|\bm{d}_{i,n}|).$ (6)
We then encode the normalized kinematic feature as
$\tilde{\bm{\phi}}_{i,n}=\mathcal{E}_{k}\!\left(\mathrm{Norm}\!\left(\left[\bm{r}_{i,n}^{\top},\tilde{\bm{d}}_{i,n}^{\top}\right]^{\top}\right)\right),$ (7)
yielding latent kinematic tokens in $\mathbb{R}^{d}$ that match the embedding dimension of the audio transformer, where $\mathcal{E}_{k}$ denotes a three-layer MLP.
Audio-Kinematic Fusion. The encoded kinematic tokens $\bm{H}_{\text{kin}}=\{\tilde{\bm{\phi}}_{i,n}\}\in\mathbb{R}^{f\times N\times d}$ are injected into the audio branch of the twin-DiT architecture to explicitly modulate audio generation with motion information. Specifically, within each audio transformer block, an auxiliary cross-attention layer is introduced after the self-attention layer, with audio latent states as queries and kinematic tokens as keys and values. We first project $\bm{H}_{\text{kin}}$ to key and value tensors. For clarity, we omit the multi-head notation:
$\bm{K}_{\text{kin}}=\mathrm{Norm}(\bm{H}_{\text{kin}}\bm{W}_{\text{kin}}^{(K)}),\quad\bm{V}_{\text{kin}}=\bm{H}_{\text{kin}}\bm{W}_{\text{kin}}^{(V)}.$ (8)
To preserve temporal alignment, we apply RoPE [39] to the temporal dimension of $\bm{K}_{\text{kin}}$, reshape $\bm{K}_{\text{kin}}$ and $\bm{V}_{\text{kin}}$ to $\mathbb{R}^{(fN)\times d}$, and apply RoPE to audio latents $\bm{H}_{a}\in\mathbb{R}^{L_{a}\times d}$, with temporal indices scaled by $f/L_{a}$, following Ovi. The kinematic-aware update $\Delta\bm{H}_{a}$ is computed as:
$\bm{Q}_{a}$ $=\mathrm{Norm}(\bm{H}_{a}\bm{W}_{a}^{(Q)}),$ (9)
$\Delta\bm{H}_{a}$ $=\mathrm{Softmax}\!\left(\frac{\bm{Q}_{a}\bm{K}_{\text{kin}}^{\top}}{\sqrt{d}}\right)\bm{V}_{\text{kin}}.$ (10)
To prevent kinematic cues from dominating semantic text conditions during early training, we further employ a learnable gating mechanism. The final hidden state $\bm{H}_{a}^{\prime}$ is updated through a residual connection modulated by a sigmoid-gated scalar $\sigma(\gamma)$, where $\gamma$ is a learnable parameter initialized to a negative value:
$\bm{H}_{a}^{\prime}=\bm{H}_{a}+\sigma(\gamma)\cdot\Delta\bm{H}_{a}.$ (11)
This design allows Tora3 to adaptively balance semantic context and motion-aware conditioning, producing audio that is both semantically appropriate and better aligned with the visual motion.
Hybrid Flow Matching
Directly reusing first-frame latents along trajectories is effective for injecting motion cues, but applying the same construction uniformly to all spatial locations can be limiting. We observe that different regions play different roles during denoising. Trajectory-conditioned regions benefit from stronger appearance anchoring to preserve motion guidance, whereas non-trajectory regions benefit from the standard flow to maintain flexible scene refinement. Motivated by this observation, we propose Hybrid Flow Matching, which modifies the probability flow inside the trajectory region $\Omega_{\text{traj}}=\left\{(i,p_{i,n},q_{i,n})\;\middle|\;i\in\{0,\dots,f-1\},\ n\in\{0,\dots,N-1\}\right\}.$ while preserving the standard flow outside it.
Outside $\Omega_{\text{traj}}$, we follow the standard Flow Matching construction on the clean video latent $x_{0}$:
$x_{t}=(1-t)x_{0}+t\epsilon,\qquad v=\epsilon-x_{0},$ (12)
where $t\in[0,1]$ denotes the flow matching timestep, $\epsilon\sim\mathcal{N}(0,I)$ denotes the standard Gaussian endpoint, and $v$ denotes the target velocity field for unconstrained regions.
Inside $\Omega_{\text{traj}}$, we replace the standard noise endpoint with the trajectory-conditioned latent $x^{\mathrm{traj}}$, yielding
$x_{t}=(1-t)x_{0}+tx^{\mathrm{traj}},\qquad v=x^{\mathrm{traj}}-x_{0}.$ (13)
Intuitively, this preserves the standard denoising behavior in unconstrained regions while anchoring trajectory-conditioned regions to the injected motion prior.
To implement this hybrid construction, we use the binary indicator mask $M\in\{0,1\}^{f\times h\times w}$ associated with $\Omega_{\text{traj}}$, defined by $M[i,p,q]=\mathbf{1}[(i,p,q)\in\Omega_{\text{traj}}]$. The resulting latent input and target velocity field are defined as
$x_{t}=(1-M)\odot\big((1-t)x_{0}+t\epsilon\big)+M\odot\big((1-t)x_{0}+tx^{\mathrm{traj}}\big),$ (14)
$v=(1-M)\odot(\epsilon-x_{0})+M\odot(x^{\mathrm{traj}}-x_{0}).$ (15)
Here, the mask $M$ is broadcast along the channel dimension to $\mathbb{R}^{f\times h\times w\times d}$. At inference time, we use the same trajectory-conditioned latent construction to initialize the flow-matching process at $t=1$, with $x_{1}=(1-M)\odot\epsilon+M\odot x^{\mathrm{traj}}$.
This construction ensures that the latent input at each spatial location is consistent with the corresponding velocity target used for supervision. In practice, it helps preserve trajectory-guided motion while reducing local inconsistencies around trajectory-conditioned regions.
During training, we keep the audio objective unchanged from Ovi and only replace its original video loss with a region-balanced objective to better balance the sparse trajectory region and its complement. To ensure smooth boundary handling, we use a soft mask $M_{\text{soft}}$ obtained by Gaussian blurring the binary mask $M$. The modified video loss is defined as
$\mathcal{L}_{\text{video}}=\lambda_{\text{out}}\mathcal{L}_{\text{out}}+\lambda_{\text{traj}}\mathcal{L}_{\text{traj}},$ (16)
where $\lambda_{\text{out}}=\lambda_{\text{traj}}=0.5$. The loss terms are defined as
$\mathcal{L}_{\text{out}}$ $=\frac{\sum\big((1-M_{\text{soft}})\odot(\hat{v}-v)^{2}\big)}{\sum(1-M_{\text{soft}})+\delta},$ (17)
$\mathcal{L}_{\text{traj}}$ $=\frac{\sum\big(M_{\text{soft}}\odot(\hat{v}-v)^{2}\big)}{\sum M_{\text{soft}}+\delta},$ (18)
where $\delta=10^{-8}$ is a small constant for numerical stability. This equal weighting prevents the sparse trajectory region from being overwhelmed by the dominant non-trajectory area, allowing the model to learn localized motion control and global appearance generation in a more balanced manner. Following Ovi, the final training objective combines the video and audio losses with weights $0.85$ and $0.15$, respectively:
$\mathcal{L}_{\text{final}}=0.85\,\mathcal{L}_{\text{video}}+0.15\,\mathcal{L}_{\text{audio}}^{\text{Ovi}},$ (19)
where $\mathcal{L}_{\text{audio}}^{\text{Ovi}}$ denotes the original audio loss used in Ovi.
Table 1: Comparison on video quality, audio quality, text alignment, motion control, and audio-video synchronization, including event-level and intensity-level motion-sound coherence metrics.
Video Quality Audio Quality Text Alignment AV Synchronization Motion Control
Method # Params AS $\uparrow$ FVD $\downarrow$ CE $\uparrow$ CU $\uparrow$ PC $\downarrow$ PQ $\uparrow$ CLAP $\uparrow$ CLIP-T $\uparrow$ FGAS $\uparrow$ ETE $\downarrow$ MAIC $\uparrow$ TE $\downarrow$
LTX-2 [17] 22.16B 4.31 989.6 3.28 6.17 2.43 6.73 0.31 0.29 0.187 0.284 0.41 -
Ovi [29] 11.66B 4.40 887.7 3.30 6.01 1.85 6.44 0.43 0.30 0.156 0.301 0.37 -
MOVA [40] 30.00B 4.63 849.8 3.05 6.31 1.77 6.95 0.46 0.31 0.201 0.236 0.49 -
AVControl [2] 22.32B 4.52 829.6 3.29 6.22 2.18 6.79 0.39 0.30 0.209 0.214 0.55 19.95
Tora3 12.25B 4.61 784.1 3.34 6.43 1.81 7.09 0.44 0.31 0.234 0.181 0.63 12.13
Experiments
Experimental Setup
Datasets. We construct PAV from filtered high-quality subsets of VGGSound [5], ACAV-100M [24], OpenVid1M [30], Pexels [32], and in-house collected data. To improve motion relevance, we use Qwen3-VL [1] to automatically filter clips whose main object exhibits basic motion patterns such as translation, rotation, sliding, parabolic motion, oscillation, scale change, or deformation. For the retained clips, we first apply SAM2 [34] to segment all objects in the first frame, and then run CoTracker3 [20] initialized from the centroid of each object to obtain trajectory annotations, resulting in 460k video clips. These annotations provide scalable kinematic supervision rather than full physical measurements. We further use Qwen3-VL-8B-Instruct and Qwen3-Omni-Captioner [54] to generate paired video and audio descriptions. For quantitative evaluation, we use 50 representative videos covering diverse motion patterns and scene types.
Metrics. We evaluate generation quality from four aspects: video quality, audio quality, text alignment, and audio-video synchronization. For video, we report Fréchet Video Distance (FVD) [42], Aesthetic Score (AS), computed as the average of MANIQA [55] and MUSIQ [21], and Trajectory Error (TE), defined as the L2 distance between the conditioning trajectories and the object tracks estimated from generated videos, following [60]. For audio, we adopt AudioBox-Aesthetics [41], including Content Enjoyment (CE), Content Usefulness (CU), Production Complexity (PC), and Production Quality (PQ). To assess text alignment, we compute CLIP [33] and CLAP [50] scores for text-to-video and text-to-audio consistency, respectively. For audio-video synchronization, we utilize FGAS in PhyAVBench [52], which computes the cosine similarity between video frames and synchronized audio features in the latent space of CAV-MAE Sync [16]. To further assess physical coherence at the motion-sound level, we introduce two additional metrics. Event Timing Error (ETE) measures the time difference between trajectory-derived events and corresponding audio onsets. Motion-Audio Intensity Correlation (MAIC) computes the Pearson correlation between a trajectory-derived motion-intensity signal and the temporally aligned audio energy envelope. We evaluate ETE on clips with discrete events and MAIC on clips with sustained motion and continuous sound evolution. Detailed computation procedures are provided in the appendix.
Implementation Details. The model is initialized from the pretrained Ovi checkpoint and trained for 30k steps on 32 NVIDIA A100 GPUs with a global batch size of 32. We use AdamW with $\beta_{1}=0.9$, $\beta_{2}=0.999$, weight decay 0.01, and learning rate $4\times 10^{-5}$. Training stability is improved with BF16 mixed precision and gradient clipping at 1.0. Kinematic features are normalized using dataset-level statistics computed from 5,000 randomly sampled training examples. We further apply trajectory condition dropout with probability $p=0.05$ to improve robustness. For the mask $M_{\text{soft}}$, the binary trajectory mask is smoothed with a Gaussian kernel of $\sigma=0.5$ to encourage smooth transitions between trajectory-conditioned regions and their surroundings. The sigmoid-gated scalar parameter $\gamma$ is initialized to 10-10.
Main Results
Table 1 compares Tora3 with four competitive baselines, LTX-2 [17], Ovi [29], MOVA [40], and AVControl [2], across video quality, audio quality, text alignment, audio-video synchronization, and motion control. AVControl is a concurrent work on motion-conditioned audio-video generation, and we include its results here for a direct comparison. Overall, Tora3 achieves the best trade-off of generation quality and motion-sound coherence. It delivers the best video realism with the lowest FVD, strong overall audio quality, and the highest FGAS, indicating improved synchronization between generated motion and sound. Tora3 also achieves the lowest ETE and the highest MAIC, showing that it more accurately aligns salient audio events with motion and contact cues, and better matches the evolution of audio intensity to motion strength. At the same time, Tora3 remains highly competitive in text alignment, matching the best CLIP-T score and achieving a strong CLAP score.
Figure 4: Qualitative comparison. Tora3 produces more stable visual motion and more coherent audio in both examples, including motion-consistent loudness variation for the moving motorcycle and better synchronized rolling and collision sounds for the rolling ball. Please refer to the supplementary video for full comparisons.
Compared with AVControl, Tora3 is particularly strong in motion-sound coherence. AVControl mainly uses trajectories as a control signal for the joint backbone, whereas Tora3 uses them as a shared kinematic prior for both modalities. Trajectory-aligned motion injection improves video motion fidelity, and kinematic-audio alignment explicitly links motion states to sound generation, leading to better FGAS, ETE, and MAIC. These gains support our central design: trajectories effectively couple visual motion and acoustic generation under a shared trajectory-grounded structure.
Figure 4 shows the qualitative advantages of Tora3 over competitive baselines. In the motorcycle example, baselines either degrade visual quality or fail to produce motion-consistent loudness variation, whereas Tora3 maintains stable visuals and more coherent audio dynamics. In the rolling-ball example, baselines struggle with rolling, collision, and post-impact motion together with synchronized sound, often producing unrealistic motion, incorrect impact timing, or audio artifacts. Tora3 models these events more faithfully with better synchronized rolling and impact sounds.
Figure 5: Effect of motion speed on generated audio. Tora3 generates audio that varies consistently with motion speed.
Figure 5 shows that Tora3 produces speed-dependent audio variations. When the box is pulled faster, the generated sound is louder and exhibits subtle timbral differences compared with the sound produced under slower motion. This highlights the advantage of using trajectories as a shared kinematic prior, allowing motion changes to be reflected more consistently in both visual dynamics and generated sound.
Ablation Studies
Shared kinematic prior across modalities. Table 2 studies whether trajectory-derived kinematic conditioning should be applied to the video branch, the audio branch, or both. Applying it only to the video branch mainly improves visual generation, yielding the best Aesthetic Score and clearly better video realism. In contrast, applying it only to the audio branch more directly improves audio quality and motion-sound coherence, as reflected by higher PQ, FGAS, and MAIC as well as lower ETE. This asymmetric behavior highlights that the two designs are complementary: trajectory injection improves motion faithfulness in the video branch, while kinematic conditioning improves event timing and intensity alignment in the audio branch. When used together, they produce the strongest overall results in video realism, audio quality, and motion-sound coherence. This confirms that trajectories are most effective as a shared kinematic prior across modalities rather than isolated controls for either branch alone.
Table 2: Effect of using trajectories as a shared kinematic prior for both modalities.
Setting AS $\uparrow$ FVD $\downarrow$ PQ $\uparrow$ FGAS $\uparrow$ ETE $\downarrow$ MAIC $\uparrow$
Neither 4.39 854.7 6.75 0.173 0.296 0.39
Video only 4.51 823.6 6.81 0.198 0.247 0.46
Audio only 4.42 845.2 6.89 0.209 0.221 0.61
Both (full) 4.47 811.8 6.93 0.225 0.193 0.66
Motion representation designs for video-branch injection. Table 3 compares three motion injection strategies for the video branch: a Tora-inspired design using an auxiliary motion encoder and fuser, a WanMove-inspired design that concatenates latent noise with trajectory-guided first-frame propagated features along the channel dimension, and our Tora3 motion representation. Tora3 consistently achieves the best overall performance, yielding the highest visual quality, the strongest audio-video synchronization, the lowest event timing error, and the lowest trajectory error, while requiring no additional parameters. This advantage likely stems from the fact that: unlike prior approaches that transform sparse control signals into separate feature spaces, Tora3 injects motion directly in the native latent space by reusing first-frame features along the prescribed trajectories. This leads to cleaner and more reliable motion control, improving trajectory following in the video branch and providing more stable motion cues to the audio branch through cross-modal interaction.
Kinematic state components. Table 4 studies the effect of progressively enriching the kinematic conditioning used by the audio branch. Starting from no kinematic conditioning, adding position and velocity already improves both synchronization and motion-sound coherence. Incorporating acceleration further strengthens performance, especially in FGAS, ETE, and MAIC. The full second-order kinematic state, which additionally includes velocity and acceleration magnitudes, achieves the best performance across all metrics. This trend is consistent with the role of different kinematic components. Position provides coarse scene context, velocity indicates how objects move over time, and acceleration is particularly informative for abrupt motion changes associated with impact-like events. Adding the magnitudes of velocity and acceleration further exposes motion strength, allowing the model to better calibrate audio intensity and temporal salience. The monotonic improvements therefore support our formulation of trajectory-derived kinematic states as a compact and effective interface between trajectories and audio generation.
Table 3: Ablation on motion representation designs for video-branch injection. Tora-style and WanMove-style denote variants inspired by Tora [60] and WanMove [8], respectively.
Method AS $\uparrow$ FGAS $\uparrow$ ETE $\downarrow$ TE $\downarrow$ # Params
Tora-style [60] 4.49 0.179 0.271 17.06 12.65B (+0.99B)
WanMove-style [8] 4.44 0.184 0.251 13.91 11.66B (+590K)
Ours 4.51 0.198 0.247 13.03 11.66B (+0)
Table 4: Effect of different kinematic state components on audio generation and motion-sound coherence.
Kinematic Signal PQ $\uparrow$ FGAS $\uparrow$ ETE $\downarrow$ MAIC $\uparrow$
None 6.75 0.173 0.296 0.39
$\bm{r}+\bm{v}$ 6.79 0.191 0.254 0.48
$\bm{r}+\bm{v}+\bm{a}$ 6.86 0.202 0.229 0.56
$\bm{r}+\bm{v}+\bm{a}+\|\bm{v}\|_{2}+\|\bm{a}\|_{2}$ 6.89 0.209 0.221 0.61
Table 5: Effect of Hybrid Flow Matching (HFM) on video quality, trajectory following, and overall AV synchronization.
Method AS $\uparrow$ FVD $\downarrow$ TE $\downarrow$ FGAS $\uparrow$
w/o HFM 4.47 811.8 12.94 0.225
w/ HFM 4.61 784.1 12.13 0.234
Hybrid flow matching. Table 5 shows that hybrid flow matching consistently improves both perceptual and distributional video quality, while also improving trajectory faithfulness and overall audio-video synchronization. In particular, it improves the Aesthetic Score from 4.47 to 4.61, reduces FVD from 811.8 to 784.1, and lowers TE from 12.94 to 12.13. These results support our region-aware hybrid construction: stronger appearance anchoring in trajectory-conditioned regions improves motion fidelity, while preserving the standard flow in non-trajectory regions helps maintain local visual coherence.
Conclusion
We presented Tora3, a trajectory-guided framework for audio-video generation that improves physical coherence by using object trajectories as a shared kinematic prior for video and audio synthesis. Tora3 integrates a trajectory-aligned motion representation, a second-order kinematics-based audio alignment module, and a hybrid flow matching objective for robust motion modeling and local consistency. We also introduced PAV, a large-scale audio-video dataset with motion-centric annotations for trajectory-aware generation. Extensive experiments show that Tora3 improves visual realism, audio quality, motion-sound coherence, and audio-video synchronization over strong open-source baselines, while remaining competitive in text alignment. Future work includes extending trajectory-level coherence to richer physical factors, such as material properties, contact dynamics, and 3D acoustic propagation.